OpenStudy (anonymous):

solve for w -4/5w-1/3=-1/2

OpenStudy (anonymous):

like i am struggling

OpenStudy (anonymous):

add (1/3) to both sides -4/5w=-1/6 now multiply both sides by five -4w=-5/6 now divide both sides by -4 w=5/24

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

not convinced of Ceeasu's answer =multiply both sides by 3 to cancel the 1/3 on the LHS -4/(5w-1)=-3/2 multiply RHS by (5w-1) and LHS by 2 -8=-15w+3 simplify w=11/15